Hello! Okay, so I'm not positive that I put this problem in the correct place, but because it was called "Finance", I assumed it to be the best option because this word problem contains coins and money sums.
The word problem:
"Jane has six more nickels than quarters and two fewer dimes than nickels. If she has a total of $2.70, how many of each type of coin does she have?"
Thank you for your help! I just wasn't sure how to solve this one because I think I'm supposed to set up a chart maybe (If you could help me with that too, that would be great!), but I'm just not sure how to go about solving it. Thank you very much!

Answer by stanbon(75887)

Jane has six more nickels than quarters and two fewer dimes than nickels. If she has a total of $2.70, how many of each type of coin does she have?"
---------------
Equations:
5n + 10d + 25q = 270 cents
------
n = q+6
d = n-2
-----
Substitute for "q" and for "d" and solve for "n":
5n + 10(n-2) + 25(n-6) = 270
5n + 10n-20 + 25n-150 = 270
40n -170 = 270
40n = 440
n = 11 (# of nickels)
---
d = n-2 = 9 (# of dimes)
---
q = n-6 = 5 (# of quarters)
================================
cheers,
Stan H.
=================